OpenStudy (anonymous):

how to write x-intercept 9, y-intercept -4 in slope intercept form

OpenStudy (aum):

The equation of a line in intercept form is: \(\Large \frac{x}{\text{x-intercept}} + \frac{y}{\text{y-intercept}} = 1\) \(\Large \frac{x}{9} + \frac{y}{-4} = 1\) Multiply throughout by 36. Put the equation in the form y = mx + b